OpenStudy (anonymous):

Please help me solve this derivative? :/ As a race car accelerated away from the starting line, its velocity at any given time is modeled by this function: V=5t^2+2t Find the rate at which the car is accelerating after 3 seconds. So supposedly the answer is 32. The derivative is 10t+2 also. My question is, how do you get the derivative? What I mean is, how do you go from V=5t^2+2t to 10t+2. If you could explain it to me you have no idea how greatfull I will be!!

OpenStudy (sleepyhead314):

do you know how derivatives work? derivative is an equation that represents the slope of an equation the slope of velocity = acceleration to get derivatives, there is something called a "Power Rule" where given x^n , the derivative would be n*x^(n-1) so for example the derivative of x^3 = 3x^2 and the derivative of 5x^2 = 5*2*x^1 = 10x do you follow so far @Isabellaxoxo ? :)